If the area og a square is 290 square fee what is the approximate length?
Found 2 solutions by Alan3354, checkley79:
Answer by Alan3354(69443)

If the area og a square is 290 square fee what is the approximate length?
----------
Area = s^2 = 290
s =~ 17.03 feet

Answer by checkley79(3341)

A=L*W
THE LENGTH & THE WIDTH ARE THE SAME
A=L^2
290=L^2
L=SQRT(290}
L=17.03 IS THE APPROXIMATE LENGTH OF THE SIDES.